A bin contains a total of 20 batteries, of which 6 defective. If you select 2 at random, without replacement, determine the probability that none of the batteries is good and at least one of the batteries is good

Verified Answer :

toporc

The probability that neither of the sample of 2 is good is:
\(P(none\ good)=\frac{14C2}{20C2}=0.479\)
The probability that at least one is good is:
P(at least one is good) = 1 - 0.479 = 0.521

Hey there, Lets solve this problem together.  

We know that a bin contains a total of 20 batteries, of which 6 defective.

Therefore


 16 good ones = 16C2

20 batteries= 20C2 

\(P = none good = 0.479 P(at least one is good) = 1 - 0.479 = 0.521\)

\(12 / 19 = probability\)